# Machine Learning Time Propagators for Time-Dependent Density Functional Theory Simulations

### v2

@@ 5 Conclusion
We show that machine learned time propagators have the potential to accelerate TDDFT calculations. The propagator generalizes well to longer time roll-out periods and higher resolution spatial grids while maintaining physical consistency and observable accuracy. Extending this work to three-dimensions would enable on-the-fly modeling of the electronic response properties of laser-excited molecules and materials in various scattering experiments that are conducted at photon sources around the globe. This would enable fast simulations that generalize well over the input parameters of the experimental setup. Rapid modeling would also enable the design of laser pulses to precisely control quantum dynamics under quantum optimal control theory [54].
@@ 5.1 Data Generation
Laser intensity ${10}^{12}\text{\,}\mathrm{-}$${10}^{14}\text{\,}\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}\mathrm{{}^{2}}$ [55, 56, 57]

### v3

@@ 5 Conclusion
We show that machine learned time propagators have the potential to accelerate TDDFT calculations for one-dimensional diatomic molecules. The propagator generalizes well to longer time roll-out periods and higher resolution spatial grids while maintaining physical consistency and observable accuracy. Extending this work to three-dimensions would enable on-the-fly modeling of the electronic response properties of laser-excited molecules and materials in various scattering experiments that are conducted at photon sources around the globe. This would enable fast simulations that generalize well over the input parameters of the experimental setup. Rapid modeling would also enable the design of laser pulses to precisely control quantum dynamics under quantum optimal control theory [werschnik_quantum_2007].
@@ 5.1 Data Generation
Laser intensity ${10}^{12}\text{\,}\mathrm{-}$${10}^{14}\text{\,}\mathrm{W}\mathrm{/}\mathrm{c}\mathrm{m}\mathrm{{}^{2}}$ [andTDDFTStudyExcitation2013, bubinElectronionDynamicsLaserassisted2011a, wozniakChapterThreeExploring2023]